OpenStudy (anonymous):

evaluate the square root of 24-x when x=6. Write the answer in simplified radical form.

OpenStudy (anonymous):

idk what radical form is

OpenStudy (janu16):

can you draw that or something

OpenStudy (anonymous):

OK so √24 is 2√3

OpenStudy (anonymous):

2√3-6

OpenStudy (anonymous):

? i got -5.33333333 idk if that correct but lemme try again

OpenStudy (anonymous):

U have to rationalize the √24

OpenStudy (anonymous):

√24=√2*√12 √12=√6 √3 √6=√2√3

OpenStudy (anonymous):

I have a pair of 2 and 3

OpenStudy (shadowlegendx):

\[\sqrt{24 - x}\] \[\sqrt{24 - 6}\]

OpenStudy (shadowlegendx):

\[\sqrt{18}\]

OpenStudy (anonymous):

oh now i get it

OpenStudy (shadowlegendx):

@usernamenotfound you are neglecting the -6

OpenStudy (anonymous):

So it's gonna be √18=√9√2 √9=√3√3

OpenStudy (anonymous):

6√2

OpenStudy (shadowlegendx):

Close \[3\sqrt{2}\]